# ReLaX-VQA: Residual Fragment and Layer Stack Extraction for Enhancing Video Quality Assessment

### v1

@@ 3.1 Spatio-temporal fragment sampling module
Frequently in VQA, researchers resize or crop frames prior to feature extraction. Resizing often introduces quality degradation as it filters or changes local textures, which are often critical in the perception of video quality, while cropping affects the alignment of global and local quality. Particularly at high spatial resolutions, these processing methods might result in a significant loss of detail. Moreover, the viewer does not only notice the overall quality but also temporal artefacts, e.g., sudden object transitions/deformations due to motion estimation errors, and/or any other spatial artefacts, e.g., blocking artefacts, colour bleeding. By extracting the key fragments with the highest dynamic changes from frame to frame, we can effectively capture both local and global information affecting video quality. Most UGC videos are short, usually containing only one shot, and redundancy in spatio-temporal information especially at high frame rates. Therefore, we propose an innovative Residual Fragment (RFs) sampling method to capture the effect of motion and spatial variations on video quality.
@@
We sampled the video at a rate of two frames every half second to obtain successive and temporally correlated frame pairs $F_{t}$ and $F_{t+1}$. We first computed the frame difference $\Delta F_{t}(x,y)$ by calculating the absolute difference of each pixel in between the extracted frame pairs to capture the inter-frame content variations, as follows:
@@
Concurrently, to better capture motion information, we applied Farneback dense optical flow estimation farneback2003two to quantify pixel motion across video frames based on temporal continuity and spatial luminance consistency. Optical flow $f_{t}(x,y)$ is computed by comparing the pixel positions between consecutive frames $F_{t}$ and $F_{t+1}$ to estimate the direction and velocity of movement of each reference pixel $(x,y)$, by solving the following equation:
@@
$\text{\small RGB}=\text{\small HSVtoRGB}\left(\left[\frac{\arctan 2(v,u)}{\pi}\times 180,\,2^{b}-1,\,\min\left(\frac{\sqrt{u^{2}+v^{2}}}{\max(|\mathbf{F}|)},1\right)\times(2^{b}-1)\right]\right)\,,$ (3)
@@
11:Find top-$K$ patches: $top\_idx\leftarrow\text{sort}(\text{argsort}(P,K))$
12:Initialise $RF\leftarrow 0$ and $P_{\text{positions}}\leftarrow[]$
13:for each index $(y,x)$ in $top\_idx$ do
14:     Extract patch: $patch\leftarrow P[y\cdot p:(y+1)\cdot p,x\cdot p:(x+1)\cdot p]$
15:     Determine new location in the target image $I_{\text{patches}}$ (size: $n\times n$):
@@
16:     Insert patch into $RF$ and update $P_{\text{positions}}\leftarrow P_{\text{positions}}\cup\{(y,x)\}$
17:end for
18:return $P,RF,P_{\text{positions}}$
@@
$S_{\Delta F_{t}}(i,j)=\sum_{x=i}^{i+p}\sum_{y=j}^{j+p}\Delta F_{t}(x,y)\quad\text{and}\quad S_{f_{t}}(i,j)=\sum_{x=i}^{i+p}\sum_{y=j}^{j+p}\|\mathbf{f}_{t}(x,y)\|\,,$ (4)
@@
where $MF(x,y)$ is the merged fragment and $\alpha=0.5$ is a weighting parameter for the two fragments, $RF_{\Delta F}$ and $RF_{f}$. The target RFs are sized at $224\times 224$ pixels to align with the model input requirements. We set the size of the top patches to $p\times p$ (where $p=16$) pixels, arranging them in a $14\times 14$ grid to achieve the target dimensions. Consequently, a total of 196 important patches need to be selected. This process is detailed in Algorithm 1(line 11). We also extracted patches from the same positions as the top frame difference patches to construct sampled spatial fragments, based on the principle of spatial position invariance. This sampling criterion is based on the hypothesis that patches with higher residual information (indicating increased movement or variation) are critical for understanding changes in video content and are more likely to convey noticeable compression artefacts. We also provided two example videos to visualise the merged fragments in Fig. 3 in the Appendix.
3.2 DNN layer stack module
@@
We observed that the lower (first) layers of CNN networks process basic image elements such as edges, colours, and textures, while deeper layers handle advanced more abstract information. Recent studieswang2023ugc ; ramsookicip2023 indicated that saliency features may be inherent in representations of different DNN layers (not in just pooled flattened features), allowing us to simulate human attention to regions of interest. Therefore in this work we used layer stacking and global pooling techniques to enhance DNN features and improve the perceptual accuracy of the prediction model. Besides this, through the integration of ViT features vaswani2017attention , which focus on global information, with ResNet-50he2016deep features, which focus on local details, we can further improve the model’s ability to process complex visual information. To support this, we conducted an ablation study to evaluate the impact of using different features alone, further validating the proposed method (see Section 4, Table 4).
@@
For each layer $n$ of ResNet-50, the feature map $m^{(n)}(x)$ possesses dimensions $C_{n}\times H_{n}\times W_{n}$, where $C_{n}$ represents the number of channels, $H_{n}$ represents the height, and $W_{n}$ represents the width. The channel mean is computed for each channel $c$ in the $n$ layer. The feature vectors $v^{(C_{n})}$ from each layer are concatenated to form a layer-stacked feature comprising all the layer features:
$v^{(n)}c=\frac{1}{H_{n}\times W_{n}}\sum{h=1}^{H_{n}}\sum_{w=1}^{W_{n}}m^{(n)}(x_{c,h,w})$ (6)
$V_{\text{CNN}}^{\text{stack}}=v^{(1)}\oplus v^{(2)}\oplus\cdots\oplus v^{(C_{l})}$ (7)
@@
$\mu=\text{mean}(m^{(n)}(x)),\quad\zeta=\text{max}(m^{(n)}(x)),\quad\sigma=\text{std}(m^{(n)}(x))$ (8)
@@
$V_{\text{ViT}}^{\text{pool}}=\mu_{\text{patch\_embedding}}\oplus\zeta_{\text{patch\_embedding}}\oplus\sigma_{\text{patch\_embedding}}$ (10)
Next, we aggregated the enhanced feature vectors from ResNet-50 and ViT to acquire DNN features for the sampled video frames. After processing all the sampled frames, we derived a feature representation of the video sequence by averaging the features of each frame $V_{\text{video}}=\frac{1}{N}\sum_{i=1}^{N}(V_{\text{CNN}}\oplus V_{\text{ViT}})$.
3.3 Quality regression module
We designed a simple and efficient MLP rosenblatt1958perceptron regression head to predict video quality scores. The model consists of three fully connected layers, with hidden layers of dimensions 256 and 128, and the last layer outputs the predicted scores. We applied the GELU activation function and a dropout rate of 0.1 after the first two layers, with the final layer directly outputting the results. Throughout the training process, we utilized the SGD optimizer and applied Cosine Annealing learning rate decayloshchilov2016sgdr . Additionally, we introduced the Stochastic Weight Averaging (SWA)izmailov2018averaging technique to enhance the model’s generalization performance. To prevent overfitting, we adopted an early stopping strategy with a patience of 5 epochs. Moreover, we incorporated batch normalization for handling the large dataset LSVQying2021patch .
3.4 Loss function
We used a composite loss function inspired by wen2021strong , where the MAE loss and Rank Loss are weighed by specific weights to achieve optimal learning. Specifically, we introduced two weighting parameters, $l1_{w}$ and $rank_{w}$, to adjust the weights of the two losses in the total loss function. MAE loss is used to calculate the average absolute difference between predicted scores and ground truth:
@@ 3.4 Loss function
where $y_{i}\in\mathbb{R}$ represents a quality score.
@@
$L_{\text{Rank}}=\frac{1}{n\cdot(n-1)}\sum_{i=1}^{n}\sum_{j=1}^{n}\max\left(0,\delta_{ij}-e(y_{\text{true},i},y_{\text{true},j})\cdot d_{ij}\right)\,,$ (12)
@@
4 Experiments
4.1 Experimental setup
UGC Datasets
Our model’s performance has been evaluated on current state-of-the-art NR-VQA datasets including CVD-2014nuutinen2016cvd2014 , KoNViD-1khosu2017konstanz , LIVE-VQCsinno2018large , and YouTube-UGCwang2019youtube . Additionally, we have conducted feature extraction and training on the large-scale LSVQying2021patch dataset and successfully migrated to smaller VQA datasets by Fine-Tuning (FT) the pre-trained model, thereby achieving better performance. We excluded 46 greyscale videos from the YouTube-UGC* dataset, leaving 1,103 videos, and 283 greyscale videos from the LSVQ** dataset, leaving 38,510 videos.
Evaluation method
We employed evaluation metrics including SRCC, KRCC, PLCC, and RMSE to assess the performance of our NR-VQA model. These metrics evaluate the monotonicity, linearity, and accuracy of the predictions. In this case, PLCC and RMSE are calculated after a nonlinear four-parameter logistic regressionseshadrinathan2010study . All results will be reported on our project GitHub page, while in Table 4.2 we only report SRCC and PLCC values.
Implementation details
We used ResNet-50he2016deep pre-trained on ImageNetkrizhevsky2012imagenet and ViT-B/16author2023visualizing pre-trained models to extract video features. The dataset was split into training, validation, and test sets in a 64%-16%-20% ratio, respectively. For the LSVQ dataset, we applied 10-fold cross-validation, allocating 9 folds for training and 1 fold for validation, to improve the generalization of our MLP model. The training was conducted over 21 iterations, with 20 epochs for the LSVQ dataset and 120 epochs for other smaller datasets. We used a batch size of 256. Specifically, for the LSVQ dataset, SGD was employed with a learning rate of 0.1 and a weight decay of 0.005, while for other datasets, the learning rate was 0.01 with a weight decay of 0.0005. SWA was implemented in the later stages of training with a learning rate of 0.05. Our loss function utilized weight parameters $l1_{w}=0.6$ and $rank_{w}=1.0$. We tested model selection with two criteria: $byrmse$, which selects the best model with the smallest RMSE in the validation set, and $bykrcc$, which selects the best model with the highest KRCC. We evaluated the performance by reporting the median and standard deviation of SRCC and PLCC across 21 iterations. All experiments utilized a workstation with an NVIDIA GeForce RTX 3090 GPU, a 6-core Intel(R) Xeon(R) W-1250 CPU, and 64.0 GB RAM. The proposed model was implemented in PyTorchpaszke2019pytorch on Python 3.8.
4.2 Performance comparison
Table 2: Performance comparison of the evaluated NR-VQA models on the four NR-VQA datasets. The red and blue entries indicate the best and second-best performance on each database for each performance metric, respectively.
Datasets CVD2014nuutinen2016cvd2014 KoNViD-1khosu2017konstanz LIVE-VQCsinno2018large YouTube-UGCwang2019youtube Overall
Models/Metrics SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
BRISQUEmittal2011blind 0.5553 0.5527 0.6781 0.6746 0.6096 0.6652 0.3517 0.3768 0.5487 0.5673
V-BLIINDS6705673 0.7306 0.7853 0.7258 0.7155 0.6925 0.6872 0.4783 0.4967 0.6568 0.6712
TLVQMkorhonen2019two 0.5399 0.5785 0.7616 0.7463 0.8133 0.7912 0.6802 0.6876 0.6988 0.7009
VIDEVALtu2021ugc 0.7663 0.8062 0.8073 0.7923 0.7725 0.7752 0.7814 0.7929 0.7819 0.7916
@@ 4.2 Performance comparison
ReLaX-VQA (ours) 0.8643 0.8895 0.8535 0.8473 0.7655 0.8079 0.8014 0.8204 0.8212 0.8413
ReLaX-VQA (w/o FT) 0.7845 0.8336 0.8312 0.8427 0.7664 0.8242 0.8104 0.8354 0.7981 0.8340
ReLaX-VQA (w/ FT) 0.8974 0.9294 0.8720 0.8668 0.8468 0.8876 0.8469 0.8652 0.8658 0.8872
We used the following datasets to train and evaluate our proposed model, CVD2014, KoNViD-1k, LIVE-VQC, YouTube-UGC and the largest dataset available, LSVQ. Based on the different strategies for training and testing we created three versions of our proposed method, ReLaX-VQA:

### v3

@@ III-A Spatio-Temporal Fragment Sampling module
Frequently in VQA, researchers resize or crop frames prior to feature extraction. Resizing often introduces quality degradation as it filters or changes local textures, which are often critical in the perception of video quality, while cropping affects the alignment of global and local quality. Particularly at high spatial resolutions, these processing methods might result in a significant loss of detail. Moreover, the viewer does not only notice the overall quality but also temporal artefacts, e.g., sudden object transitions/deformations due to motion estimation errors, and/or any other spatial artefacts, e.g., blocking artefacts, colour bleeding. By extracting the key fragments with the highest dynamic changes from frame to frame, we can effectively capture both local and global information affecting video quality. Most UGC videos are short, usually containing only one shot and redundancy in spatio-temporal information, especially at high frame rates. Therefore, we propose an innovative Residual Fragment (RFs) sampling method to capture the effect of temporal and spatial variations on video quality.
@@
We sampled the video at a rate of two frames every half-second to obtain successive and temporally correlated frame pairs $F_{t}$ and $F_{t+1}$. We first computed the frame difference $\Delta F_{t}(x,y)$ by calculating the absolute difference of each pixel in between the extracted frame pairs to capture the inter-frame content variations, as follows:
@@
Concurrently, to better capture motion information, we applied Farneback dense optical flow estimation [45] to quantify pixel displacement across video frames based on temporal continuity and spatial luminance consistency. Optical flow $f_{t}(x,y)$ is computed by comparing the pixel positions between consecutive frames $F_{t}$ and $F_{t+1}$ to estimate the direction and velocity of movement of each reference pixel $(x,y)$, by solving the following equation:
@@
$\text{\small RGB}=\text{\small HSVtoRGB}\left(\left[\frac{\arctan 2(v,u)}{\pi}\times 180,\,2^{b}-1,\right.\right.\\ \left.\left.\min\left(\frac{\sqrt{u^{2}+v^{2}}}{\max(|\mathbf{F}|)},1\right)\times(2^{b}-1)\right]\right),$ (3)
@@
11:end for
12:Find top-$K$ patches: $top\_idx\leftarrow\text{sort}(\text{argsort}(P,K))$
13:Initialise $RF\leftarrow 0$ and $P_{\text{positions}}\leftarrow[]$
14:for each index $(y,x)$ in $top\_idx$ do
15:     Extract patch:
16:     $patch\leftarrow P[y\cdot p:(y+1)\cdot p,x\cdot p:(x+1)\cdot p]$
17:     Determine new location in the target image $I_{\text{patches}}$ (size: $n\times n$):
@@
18:     Insert patch into $RF$
19:     and update $P_{\text{positions}}\leftarrow P_{\text{positions}}\cup\{(y,x)\}$
20:end for
21:return $P,RF,P_{\text{positions}}$
@@
$S_{\Delta F_{t}}(i,j)=\sum_{x=i}^{i+p}\sum_{y=j}^{j+p}\Delta F_{t}(x,y),$ (4)
$S_{f_{t}}(i,j)=\sum_{x=i}^{i+p}\sum_{y=j}^{j+p}\|\mathbf{f}_{t}(x,y)\|,$ (5)
@@
where $MF(x,y)$ is the merged fragment and $\alpha=0.5$ is a weighting parameter for the two fragments, $RF_{\Delta F}$ and $RF_{f}$. The target RFs are sized at $224\times 224$ pixels to align with the model input requirements. We set the size of the top patches to $p\times p$ (where $p=16$) pixels, arranging them in a $14\times 14$ grid to achieve the target dimensions. Consequently, a total of 196 important patches need to be selected. This process is detailed in Algorithm 1(line 11). We also extracted patches from the same positions as the top frame difference patches to construct sampled spatial fragments based on the principle of spatial position invariance. This sampling criterion is based on the hypothesis that patches with higher residual information (indicating increased movement or variation) are critical for understanding changes in video content and are more likely to convey noticeable compression artefacts.
We also provide frames from two example videos to visualise the merged fragments in Fig. 3.
Figure 3: Illustration of examples sampled from the YouTube-UGC [3] dataset. Top-row: $TelevisionClip\_1080P-68c6.mkv$. Bottom-row: $Sports\_2160P-0455.mkv$. GT refers to ground truth quality scores. Shown here are the scores predicted using ReLaX-VQA (w/o FT).
Here, we present the sampled frames and their corresponding ViT attention visualisations, along with the merged fragments (ours) and their ViT attention. These visualisations highlight the attention focus of the model. As shown in Fig. 3, the first video, $TelevisionClip\_1080P-68c6.mkv$, is a clip showing a man wearing an Angry Birds print moving in a snowy environment. The second video $Sports\_2160P-0455.mkv$ is a scene from a camera positioned slightly above the dashboard of a racing car on a track, providing a clear view through the windshield. The scene shows changes in perspective, including significant jittering. The safety barriers on the curve appear moving faster. The videos can be found in the YouTube-UGC dataset [3]. It can be observed that we get good results even when using our frozen ReLaX-VQA model without fine-tuning (ReLaX-VQA (w/o FT) to predict video quality scores.
III-B DNN Feature Extraction module
@@
We observed that the lower (first) layers of CNN networks process basic image elements such as edges, colours, and textures, while deeper layers handle more advanced abstract information. Recent studies [46, 47] indicated that saliency features may be inherent in representations of different DNN layers (not in just pooled flattened features), allowing us to simulate human attention to regions of interest. Therefore, in this work, we used layer stacking and global pooling techniques to enhance DNN features and improve the perceptual accuracy of the prediction model. Besides this, through the integration of ViT features [21], which focus on global information, with ResNet-50 [20] features, which focus on local details, we can further improve the model’s ability to process complex visual information. To support this, we conducted an ablation study to evaluate the impact of using different features alone, further validating the proposed method (see Section IV, Table V).
@@
For each layer $n$ of ResNet-50, the feature map $m^{(n)}(x)$ has dimensions $C_{n}\times H_{n}\times W_{n}$, where $C_{n}$ represents the number of channels, $H_{n}$ represents the height, and $W_{n}$ represents the width. The channel mean is computed for each channel $c$ in the $n$ layer. The feature vectors $v^{(C_{n})}$ from each layer are concatenated to form a layer-stacked feature comprising all the layer features:
$v^{(n)}c=\frac{1}{H_{n}\times W_{n}}\sum{h=1}^{H_{n}}\sum_{w=1}^{W_{n}}m^{(n)}(x_{c,h,w}),$ (7)
$V_{\text{CNN}}^{\text{stack}}=v^{(1)}\oplus v^{(2)}\oplus\cdots\oplus v^{(C_{l})}.$ (8)
@@
$\mu=\text{mean}(m^{(n)}(x)),\zeta=\text{max}(m^{(n)}(x)),\sigma=\text{std}(m^{(n)}(x)).$ (9)
@@
$V_{\text{ViT}}^{\text{pool}}=\mu_{\text{patch\_emb}}\oplus\zeta_{\text{patch\_emb}}\oplus\sigma_{\text{patch\_emb}}$ (11)
Next, we aggregated the enhanced feature vectors from ResNet-50 and ViT to acquire DNN features for the sampled video frames. After processing all the sampled frames, we derived a feature representation of the video sequence by averaging the features of each frame:
$V_{\text{video}}=\frac{1}{N}\sum_{i=1}^{N}(V_{\text{CNN}}\oplus V_{\text{ViT}}).$ (12)
III-C Quality Regression module
We designed a simple and efficient MLP [48] regression head to predict video quality scores. The model consists of three fully connected layers, with hidden layers of dimensions 256 and 128, and the last layer outputs the predicted scores. We applied the GELU activation function and a dropout rate of 0.1 after the first two layers, with the final layer directly outputting the results. Throughout the training process, we utilized the SGD optimizer and applied Cosine Annealing learning rate decay [49]. Additionally, we introduced the Stochastic Weight Averaging (SWA) [50] technique to enhance the model’s generalisation performance. To prevent overfitting, we adopted an early stopping strategy with a patience of 5 epochs. Moreover, we incorporated batch normalization for handling the large dataset LSVQ [9].
III-D Loss function
We used a composite loss function inspired by  [51], where the MAE loss and Rank Loss are weighed by specific weights to achieve optimal learning. Specifically, we introduced two weighting parameters, $l1_{w}$ and $rank_{w}$, to adjust the weights of the two losses in the total loss function. MAE loss is used to calculate the average absolute difference between predicted scores and ground truth:
@@ III-D Loss function
where $y_{i}\in\mathbb{R}$ represents a quality score, $n$ is the number of videos.
@@
$L_{\text{Rank}}^{\text{ij}}=\max\left(0,\delta_{ij}-e(y_{\text{true},i},y_{\text{true},j})\cdot d_{ij}\right),$ (14)
$L_{\text{Rank}}=\frac{1}{n\cdot(n-1)}\sum_{i=1}^{n}\sum_{j=1}^{n}L_{\text{Rank}}^{\text{ij}},$ (15)
@@
TABLE II: Performance comparison of the evaluated NR-VQA models on the four NR-VQA datasets. The Red, blue, and boldface entries indicate the 1st, 2nd and 3rd best performance on each database for each performance metric, respectively. The $\dagger$ denotes the exclusion of 4K videos.
Datasets CVD2014 [22] KoNViD-1k [23]
Models/Metrics SRCC KRCC PLCC RMSE SRCC KRCC PLCC RMSE
BRISQUE [30] 0.5553 0.3895 0.5527 18.4752 0.6781 0.4935 0.6746 0.4753
V-BLIINDS [31] 0.7306 0.5393 0.7853 13.7267 0.7258 0.5322 0.7155 0.4498
TLVQM [33] 0.5399 0.4006 0.5785 18.0832 0.7616 0.5635 0.7463 0.4285
VIDEVAL [34] 0.7663 0.5634 0.8062 13.1151 0.8073 0.6036 0.7923 0.3928
@@
Fast-VQA-B (w/o FT) [14] 0.8124 0.6125 0.8221 12.7920 0.8581 0.6684 0.8551 0.3450
Fast-VQA-B (w/ FT) [14] 0.8341 0.6466 0.8456 11.6044 0.8869 0.7052 0.8776 0.3131
DOVER (Technical Branch) [44] 0.8571 0.6966 0.8588 11.0493 0.8736 0.6962 0.8852 0.2920
ReLaX-VQA (ours) 0.8643 0.6960 0.8895 9.8185 0.8535 0.6594 0.8473 0.3370
ReLaX-VQA (w/o FT) 0.7845 0.5930 0.8336 12.2445 0.8312 0.6418 0.8427 0.3466
ReLaX-VQA (w/ FT) 0.8974 0.7299 0.9294 8.1812 0.8720 0.6881 0.8668 0.3211
Datasets LIVE-VQC [24] YouTube-UGC [3]
Models/Metrics SRCC KRCC PLCC RMSE SRCC KRCC PLCC RMSE
BRISQUE [30] 0.6096 0.4420 0.6652 12.7480 0.3517 0.2416 0.3768 0.6349
V-BLIINDS [31] 0.6925 0.5046 0.6872 12.4031 0.4783 0.3396 0.4967 0.5949
TLVQM [33] 0.8133 0.6231 0.7912 10.4409 0.6802 0.4892 0.6876 0.4977
VIDEVAL [34] 0.7725 0.5874 0.7752 10.7846 0.7814 0.5906 0.7929 0.4177
[…]
Fast-VQA-B (w/o FT) [14] 0.8261 0.6344 0.8466 9.4487 0.7241 0.5266 0.7364 0.4728
Fast-VQA-B (w/ FT) [14] 0.8484 0.6658 0.8703 8.8464 0.8385 0.6444 0.8302 0.3880
DOVER (Technical Branch) [44] 0.7898 0.6098 0.8573 9.0036 0.8181 0.6290 0.8197 0.3559
ReLaX-VQA (ours) 0.7655 0.5785 0.8079 9.8596 0.8014 0.6167 0.8204 0.3801
ReLaX-VQA (w/o FT) 0.7664 0.5812 0.8242 9.8201 0.8104 0.6131 0.8354 0.3768
ReLaX-VQA (w/ FT) 0.8468 0.6649 0.8876 7.9869 0.8469 0.6623 0.8652 0.3437
IV Experiments
IV-A Experimental setup
UGC Datasets
Our model’s performance has been evaluated on current state-of-the-art NR-VQA datasets including CVD-2014 [22], KoNViD-1k [23], LIVE-VQC [24], and YouTube-UGC [3]. Additionally, we have conducted feature extraction and training on the large-scale LSVQ [9] dataset and successfully migrated to smaller VQA datasets by Fine-Tuning (FT) the pre-trained model, thereby achieving better performance. We excluded 46 greyscale videos from the YouTube-UGC* dataset, leaving 1,103 videos, and 283 greyscale videos from the LSVQ** dataset, leaving 38,510 videos.
Evaluation method
We employed various common statistical evaluation metrics, including SRCC, KRCC, PLCC, and RMSE, to assess the performance of our NR-VQA model. These metrics evaluate the monotonicity, linearity, and accuracy of the predictions. In this case, PLCC and RMSE are calculated after a nonlinear four-parameter logistic regression [52]. All results will be reported on our project GitHub page, while in Table III-D we only report SRCC and PLCC values as below, where $\beta$ parameters are fitted using least squares
$f(x)=\beta_{2}+\frac{\beta_{1}-\beta_{2}}{1+e^{-x+\frac{\beta_{3}}{|\beta_{4}|}}}.$ (18)
Implementation details
We used ResNet-50 [20] pre-trained on ImageNet [53] and ViT-B/16 [54] pre-trained models to extract video features. The dataset was split into training and test sets in an 80%-20% ratio, respectively. For the LSVQ dataset, we applied 10-fold cross-validation, allocating 9 folds for training and 1 fold for validation, to improve the generalization of our MLP model. The training was conducted over 21 iterations, with 20 epochs for the LSVQ dataset and 120 epochs for other smaller datasets. We used a batch size of 256. Specifically, for the LSVQ dataset, SGD was employed with a learning rate of 0.1 and a weight decay of 0.005, while for other datasets, the learning rate was 0.01 with a weight decay of 0.0005. SWA was implemented in the later stages of training with a learning rate of 0.05. Our loss function utilized weight parameters $l1_{w}=0.6$ and $rank_{w}=1.0$. We tested model selection with two criteria: $byrmse$, which selects the best model with the smallest RMSE in the validation set, and $bykrcc$, which selects the best model with the highest KRCC. We evaluated the performance by reporting the median and standard deviation of SRCC and PLCC across 21 iterations. All experiments utilized a workstation with an NVIDIA GeForce RTX 3090 GPU, a 6-core Intel(R) Xeon(R) W-1250 CPU, and 64.0 GB RAM. The proposed model was implemented in PyTorch [55] on Python 3.8.
IV-B Performance comparison
We used the following datasets to train and evaluate our proposed model, CVD2014, KoNViD-1k, LIVE-VQC, YouTube-UGC and the largest dataset available, LSVQ. Based on the different strategies for training and testing, we created three versions of our proposed method, ReLaX-VQA: